Give exact and approximate solutions to three decimal places.
y^2 - 10y + 25 = 36

Answer by scott8148(6628)

factoring ___ (y - 5)^2 = 6^2

taking square root ___ y - 5 = ±6

y - 5 = 6 ___ y = 11

y - 5 = -6 ___ y = -1